OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

@ganeshie8 @jigglypuff314

OpenStudy (anonymous):

@hartnn @Hero @ParthKohli

jigglypuff314 (jigglypuff314):

|a - b| = c take out the absolute values and get a-b = c and a-b = -c

OpenStudy (anonymous):

x - 0.525 = -0.01

jigglypuff314 (jigglypuff314):

yes and solve for x and don't forget that you would get two answers

OpenStudy (anonymous):

0.525 is what came up..

jigglypuff314 (jigglypuff314):

x - 0.525 = -0.01 +0.525 +0.525 x = 0.525 - 0.01 = ?

OpenStudy (anonymous):

0.515